OpenStudy (anonymous):

Ten dice are rolled. What is the probability that exactly one die shows a one.

OpenStudy (anonymous):

exactly one?

OpenStudy (anonymous):

Exactly one

OpenStudy (anonymous):

binomial for this one one "one" and 9 "not ones" \[\binom{10}{1}\left(\frac{1}{6}\right)\times \left(\frac{5}{6}\right)^9\]

OpenStudy (anonymous):

Thanks, I had a feeling it would be that ;)

OpenStudy (stamp):

1/6 chance that one of the ten dice is a one. The other five dice each have a 5/6 chance of being something other than one. 1/6 * nine instances of *(5/6)